# Gibbs-preserving operations requiring infinite amount of quantum coherence

### v2

@@ Abstract
Preliminaries.— We begin by introducing relevant settings and frameworks. (See Appendix A for more extensive descriptions.) Throughout this work, we consider a situation where systems are surrounded by a thermal bath with an arbitrary finite inverse temperature $\beta$. We assume that the specification of a system $X$ always comes with its Hamiltonian $H_{X}=\sum_{i=1}^{d_{X}}E_{X,i}|{i}\rangle\!\langle{i}|$ with dimension $d_{X}$. Then, the thermal Gibbs state in system $X$ is written by $\tau_{X}=e^{-\beta H_{X}}/\mathrm{Tr}(e^{\beta H_{X}})$.
We consider a quantum channel, i.e., completely-positive trace-preserving (CPTP) map, from a system $S$ to another system $S^{\prime}$. A central class of quantum channels we consider is the set of Gibbs-preserving Operations. As the name suggests, these are the operations that map Gibbs states to Gibbs states. Here, we employ a generalized notion of Gibbs-preserving Operations, in which input and output systems can generally be different Janzing et al. (2000); Renes (2014); Faist et al. (2019, 2021); Sagawa et al. (2021). Namely, we call a channel $\Lambda:S\to S^{\prime}$ Gibbs-preserving if $\Lambda(\tau_{S})=\tau_{S^{\prime}}$.
@@
where $H_{\rm tot}=H_{S}\otimes\mathds{1}_{E}+\mathds{1}_{S}\otimes H_{E}=H_{S^{\prime}}\otimes\mathds{1}_{E^{\prime}}+\mathds{1}_{S^{\prime}}\otimes H_{E^{\prime}}$ is the total Hamiltionian Janzing et al. (2000); Horodecki et al. (2009).
@@
where $\{\lambda_{i}\}_{i}$ and $\{|e_{i}\rangle\}_{i}$ are the sets of eigenvalues and eigenstates of a state $\rho$ such that $\rho=\sum_{i}\lambda_{i}|{e_{i}}\rangle\!\langle{e_{i}}|$. Quantum Fisher information is a well-known coherence quantifier that comes with a natural operational interpretation Yadin and Vedral (2016); Marvian (2022).
@@
The key idea in evaluating this is to connect our setting to the recent trade-off relation between coherence cost for channel implementation and the degree of reversibility of the channel to implement Tajima et al. (2022), which was shown to unify, e.g., the Wigner-Araki-Yanase theorems on quantum processes Wigner (1952); Araki and Yanase (1960); Ozawa (2002a); Tajima and Nagaoka (2019); Kuramochi and Tajima (2023); Emori and Tajima (2023); Ozawa (2002b); Tajima et al. (2018, 2020); Tajima and Saito (2021) and the Eastin-Knill theorems on quantum error correcting codes Eastin and Knill (2009); Faist et al. (2020); Kubica and Demkowicz-Dobrzański (2021); Zhou et al. (2021); Yang et al. (2022); Tajima and Saito (2021)—see Appendix A for details. In light of this, we find that the following class of Gibbs-preserving Operations plays a central role.
@@
The quantity $H_{S}-\Lambda^{\dagger}(H_{S^{\prime}})$ is an operator that corresponds to the local energy change in the system, and the forms in (6) and (7) indicate that ${\cal C}(\Lambda,\mathbb{P})$ measures the off-diagonal element of this operator with respect to the reversible states. More discussions about this quantity can be found in Ref. Tajima et al. (2022).
We are now in the position to present our first main result, which establishes a universal lower bound for the coherence cost for pairwise reversible Gibbs-preserving Operations. (Proof in Appendix B.)
@@
This particularly establishes a demanding coherence cost in the small error regime. Notably, Theorem 2 implies that no pairwise reversible Gibbs-preserving Operation $\Lambda$ with ${\cal C}(\Lambda,\mathbb{P})>0$ can be exactly implemented with a finite amount of coherence cost, as the lower bound diverges as $\epsilon$ approaches 0.
Therefore, the problem of whether cost-diverging Gibbs-preserving Operations exists reduces to whether there exists a pairwise reversible Gibbs-preserving Operation $\Lambda$ and a reversible pair $\mathbb{P}$ such that ${\cal C}(\Lambda,\mathbb{P})>0$ at all. The following result not only shows the existence of such operations but provides a continuous family of those.
@@
We prove this in Appendix C, which also provides an explicit construction of the corresponding pairwise reversible Gibbs-preserving Operation. Here is an illustrative example encompassed in Theorem 3. Let $S$ and $S^{\prime}$ be qubit systems with $H_{S}=|{1}\rangle\!\langle{1}|$ and $H_{S^{\prime}}=0$. Since $\tau_{S,0}=1/(1+e^{-\beta})$, $\tau_{S,1}=e^{-\beta}/(1+e^{-\beta})$, and $\tau_{S^{\prime},0}=\tau_{S^{\prime},1}=1/2$, these systems satisfy (9) for arbitrary finite temperature. The corresponding pairwise reversible Gibbs-preserving Operation $\Lambda:S\to S^{\prime}$ is
@@
It is insightful to see the structural difference between the Gibbs-preserving Operations in Eqs. (2) and (10). The one in (2) can create coherence from an incoherent state input state $|{1}\rangle\!\langle{1}|$. On the other hand, the channel in (10) cannot create coherence at all—in fact, output states are always incoherent for any input states. Instead, it can perform a measurement in a coherent basis. As we show in Appendix E, the coherent cost for the channel in (2) is upper bounded by ${\cal F}(\eta)+{\cal F}(\sigma)$, which corresponds to the sum of coherence that can be created by the channel. This shows a drastic asymmetry between creation and detection of coherence when it comes to its realization.
@@
Theorems 2 and 3 provide an overview of the classification of Gibbs-preserving Operations (Fig. 1). We remark that not all pairwise reversible Gibbs-preserving Operations come with a diverging coherence cost (e.g., identity channel)—Theorem 2 ensures the infinite cost only when there is a reversible pair $\mathbb{P}$ satisfying ${\cal C}(\Lambda,\mathbb{P})>0$. On the other hand, our results do not rule out the possibility that all cost-diverging Gibbs-preserving Operations are pairwise reversible.
@@
We prove this in Appendix D, which also provides an explicit expression of the pair $(\rho,\sigma)$ of states. Theorem 4 shows that when the available coherence resource is restricted to be finite, the range of transformations achievable by Gibbs-preserving Operations is strictly narrowed. We stress that the condition (9) is extremely mild—for virtually all scenarios Theorem 4 puts nontrivial restrictions on the physically available state transformations by Gibbs-preserving Operations.
@@
Proof can be found in Appendix F. This particularly ensures that the quantity ${\cal C}(\Lambda,\mathbb{P})$ defined in (6) serves as a key quantity that characterizes the coherence cost for a certain class of Gibbs-preserving Operations, which in turn provides an operational interpretation to this quantity.
Theorem 5 ensures the existence of a Gibbs-preserving Operation that almost achieves the lower bound. However, it does not tell much about the general upper bound that could be applied to an arbitrary Gibbs-preserving Operation. In the following, we show that, by giving up obtaining the form that almost matches the lower bound, we can obtain the general sufficient coherence cost that can be universally applied to all Gibbs-preserving Operations. In fact, we find that the applicability of our bound is much beyond Gibbs-preserving Operations—it gives a sufficient coherence cost for an arbitrary quantum channel. (Proof in Appendix F.)
@@
Let us now remark a unique characteristic of our results in relation to the previous result for trivial Hamiltonian. When input and output states are only equipped with trivial Hamiltonian, every state becomes an incoherent state, i.e., invariant under Hamiltonian evolution, and thus coherence loses its status as a precious resource. Therefore, the meaningful question in such a setting is to ask the work cost (the minimum number of work bit required) for implementing unital channels using Noisy Operations Horodecki et al. (2003), which respectively corresponds to the Gibbs-preserving and Thermal Operations for trivial Hamiltonian. Ref. Faist et al. (2015b) showed that the minimum work cost is upper bounded by a quantity scaling as $\sim\log(1/\epsilon)$ with implementation error $\epsilon$, which particularly diverges at the limit of exact implementation. Nevertheless, its lower bound has still not been established, and therefore it is still unclear if this diverging cost is a fundamental phenomenon or merely an artifact of their specific construction, which is based on a decoupling technique Rio et al. (2011). On the other hand, our Theorem 2 provides a lower bound that scale with $1/\epsilon$, which is complemented by upper bounds in Theorems 5 and 6 with the same scaling. To the best of our knowledge, our results are the first ones that establish the inherently diverging thermodynamic cost for implementing Gibbs-preserving Operations.
@@
Acknowledgments.— We thank Kaito Watanabe and Bartosz Regula for fruitful discussions. H.T. was supported by JSPS Grants-in-Aid for Scientific Research No. JP19K14610 and No. JP22H05250, and MEXT KAKENHI Grant-in-Aid for Transformative Research Areas B “Quantum Energy Innovation” Grant Numbers 24H00830 and 24H00831, and JST PRESTO No. JPMJPR2014, JST MOONSHOT No. JPMJMS2061. R.T. acknowledges the support of JSPS KAKENHI Grant Number JP23K19028, JP24K16975, JST, CREST Grant Number JPMJCR23I3, Japan, and MEXT KAKENHI Grant-in-Aid for Transformative Research Areas A “Extreme Universe” Grant Number JP24H00943.
Appendix A Background and setting
A.1 Quantum resource theories
@@
$\Lambda(\sigma)\in\mathbb{F},\ \forall\Lambda\in\mathbb{O}$ (14)
for a set $\mathbb{F}$ of free states and a set $\mathbb{O}$ of free operations. Employing the flexibility in choosing different sets of free operations is usually effective when $\mathbb{O}$ comes with a complicated structure and is difficult to analyze. For instance, in the case of entanglement, it is notoriously hard to study the full potential of LOCC, and therefore several classes of other operations, which include LOCC as their subset, were investigated. One such set is separability-preserving operations, which is the maximal set that satisfies (14). This admits a great simplification of the analysis and results in significant insights into entanglement transformation Brandão and Plenio (2010); Brandao and Datta (2011); Regula et al. (2019).
@@
A.2 Thermal and Gibbs-preserving Operations
@@
$\Lambda(\rho)=\mathrm{Tr}_{E^{\prime}}\left(U\rho\otimes\tau_{S}U^{\dagger}\right),\quad[U,H_{\rm tot}]=0$ (16)
where $H_{\rm tot}=H_{S}\otimes\mathds{1}_{E}+\mathds{1}_{S}\otimes H_{E}=H_{S^{\prime}}\otimes\mathds{1}_{E^{\prime}}+\mathds{1}_{S^{\prime}}\otimes H_{E^{\prime}}$ is the total Hamiltonian. On the other hand, one can also consider a broader class based on the axiomatic formulation, which is only restricted by the minimum requirement that they should map Gibbs states to Gibbs states. This is known as Gibbs-preserving Operations—a channel $\Lambda:S\to S^{\prime}$ is called a Gibbs-preserving operation if
$\Lambda(\tau_{S})=\tau_{S^{\prime}}.$ (17)
@@
A.3 Quantum coherence in energy eigenbasis
@@
$\Pi_{E}=\sum_{n:E_{n}=E}|{n}\rangle\!\langle{n}|.$ (18)
Then, a state $\rho$ has energetic coherence if $\sum_{E}\Pi_{E}\rho\Pi_{E}\neq\rho$. Equivalently, a state $\rho$ has nonzero coherence if $e^{-Ht}\rho e^{iHt}\neq\rho$ for some $t\in\mathbb{R}$.
The letter expression particularly allows us to formalize energetic coherence in relation to a group action represented by a unitary representation of ${\rm U}(1)$ (or $\mathbb{R}$ if the Hamiltonian contains relatively irrational eigenvalues). Namely, incoherent states—states that do not have coherence—are equivalent to the states invariant under a unitary representation $\{e^{-iHt}\}_{t}$. This observation provides a way of quantifying the amount of energetic coherence employing the resource theory of asymmetry Gour and Spekkens (2008); Marvian Mashhad (2012), which considers states invariant under action of a unitary representation of a group $G$ as free states, i.e., $\mathbb{F}=\{\sigma\,|\,U_{g}\sigma U_{g}^{\dagger}=\sigma,\ \forall g\in G\}$, and the operations covariant with such group actions as free operations, i.e., $\mathbb{O}=\{\Lambda:S\to S^{\prime}\,|\,U_{S^{\prime},g}\Lambda(\rho)U_{S^{\prime},g}^{\dagger}=\Lambda(U_{S,g}\rho U_{S,g}^{\dagger}),\ \forall g\in G\}$ where $U_{X,g}$ is a unitary representation acting on a system $X$. One can then realize that incoherent states coincide with the set of free states in the framework of resource theory of asymmetry with U(1) group with the representation $U_{t}=e^{-iHt}$, equipped with
$\mathbb{F}_{\rm inc}=\{\sigma\,|\,e^{-iHt}\sigma e^{iHt}=\sigma,\ \forall t\}$ (19)
@@
$\mathbb{O}_{\rm cov}=\{\Lambda:S\to S^{\prime}\,|\,e^{-iH_{S^{\prime}}t}\Lambda(\rho)e^{iH_{S^{\prime}}t}=\Lambda(e^{-iH_{S}t}\rho e^{iH_{S}t})\ \forall t\}.$ (20)
@@
$\Lambda(\rho)=\mathrm{Tr}_{E^{\prime}}\left(U\rho\otimes\sigma U^{\dagger}\right),\quad[U,H_{S}\otimes\mathds{1}_{E}+\mathds{1}_{S}\otimes H_{E}]=0,\ \sigma\in\mathbb{F}_{\rm inc}$ (21)
for some systems $E$ and $E^{\prime}$. The similarity between (16) and (21) represents an intriguing interplay between resource theories for quantum thermodynamics and energetic coherence. Indeed, since Gibbs states do not have energetic coherence by definition, we immediately notice that
$\mathbb{O}_{\rm TO}\subseteq\mathbb{O}_{\rm cov}.$ (22)
@@
A.4 Coherence cost for Gibbs-preserving operations

### v3

@@ Abstract
Preliminaries.— We begin by introducing relevant settings and frameworks. (See Sec. I of the Supplemental Material  [^1] See Supplemental Material for details of backgrounds, proofs, and extended discussions, which includes Refs. Gour and Spekkens (2008); Chitambar and Gour (2019); Brandão and Plenio (2010); Brandao and Datta (2011); Regula et al. (2019); Yunger Halpern et al. (2016); Brandão et al. (2015); Marvian Mashhad (2012); Kudo and Tajima (2023); Yamaguchi and Tajima (2023a, b); Shitara and Tajima (2023); Keyl and Werner (1999); Hayden and Preskill (2007); Datta (2009); Regula (2017); Rockafellar (2015); Wilming (2022); Shiraishi and Takagi (2024) for more extensive descriptions.) Throughout this work, we consider a situation where systems are surrounded by a thermal bath with an arbitrary finite inverse temperature $\beta$. We assume that the specification of a system $X$ always comes with its Hamiltonian $H_{X}=\sum_{i=1}^{d_{X}}E_{X,i}|{i}\rangle\!\langle{i}|$ with dimension $d_{X}$. Then, the thermal Gibbs state in system $X$ is written by $\tau_{X}=e^{-\beta H_{X}}/\mathrm{Tr}(e^{\beta H_{X}})$.
We consider a quantum channel, i.e., completely-positive trace-preserving (CPTP) map, from a finite-dimensional system $S$ to another finite-dimensional system $S^{\prime}$. A central class of quantum channels we consider is the set of Gibbs-preserving Operations. As the name suggests, these are the operations that map Gibbs states to Gibbs states. Here, we employ a generalized notion of Gibbs-preserving Operations, in which input and output systems can generally be different Janzing et al. (2000); Renes (2014); Faist et al. (2019, 2021); Sagawa et al. (2021). Namely, we call a channel $\Lambda:S\to S^{\prime}$ Gibbs-preserving if $\Lambda(\tau_{S})=\tau_{S^{\prime}}$.
@@
where $H_{\rm tot}=H_{S}\otimes\mathds{1}_{E}+\mathds{1}_{S}\otimes H_{E}=H_{S^{\prime}}\otimes\mathds{1}_{E^{\prime}}+\mathds{1}_{S^{\prime}}\otimes H_{E^{\prime}}$ is the total Hamiltionian Janzing et al. (2000); Horodecki et al. (2009); foo .
@@
where $\{\lambda_{i}\}_{i}$ and $\{|e_{i}\rangle\}_{i}$ are the sets of eigenvalues and eigenstates of a state $\rho$ such that $\rho=\sum_{i}\lambda_{i}|{e_{i}}\rangle\!\langle{e_{i}}|$. Quantum Fisher information is a well-known coherence quantifier that comes with a natural operational interpretation Yadin and Vedral (2016); Marvian (2022); Zhang et al. (2017); Takagi (2019).
@@
The key idea in evaluating this is to connect our setting to the recent trade-off relation between coherence cost for channel implementation and the degree of reversibility of the channel to implement Tajima et al. (2022), which was shown to unify, e.g., the Wigner-Araki-Yanase theorems on quantum processes Wigner (1952); Araki and Yanase (1960); Ozawa (2002a); Tajima and Nagaoka (2019); Kuramochi and Tajima (2023); Emori and Tajima (2023); Ozawa (2002b); Tajima et al. (2018, 2020); Tajima and Saito (2021) and the Eastin-Knill theorems on quantum error correcting codes Eastin and Knill (2009); Faist et al. (2020); Kubica and Demkowicz-Dobrzański (2021); Zhou et al. (2021); Yang et al. (2022); Tajima and Saito (2021)—see Sec. I of the Supplemental Material for details. In light of this, we find that the following class of Gibbs-preserving Operations plays a central role.
@@
The quantity $H_{S}-\Lambda^{\dagger}(H_{S^{\prime}})$ is an operator that corresponds to the local energy change in the system, and the forms in (6) and (7) indicate that ${\cal C}(\Lambda,\mathbb{P})$ measures the off-diagonal element of this operator with respect to the reversible states. More discussions about this quantity can be found in Sec. I.5 of the Supplemental Material (see also Ref. Tajima et al. (2022)).
We are now in the position to present our first main result, which establishes a universal lower bound for the coherence cost for pairwise reversible Gibbs-preserving Operations. (Proof in Sec. II of the Supplemental Material.)
@@
This particularly establishes a demanding coherence cost in the small error regime. In Sec. II of the Supplemental Material, we extend Theorem 1 to the cases without perfect pairwise reversibility, as well as to the form that does not directly depend on the maximum range $\Delta(H_{S})$ or $\Delta(H_{S^{\prime}})$ of the Hamiltonians, showing the potential of obtaining a similar lower bound applicable to unbounded Hamiltonians in infinite-dimensional systems.
Theorem 1 implies that no pairwise reversible Gibbs-preserving Operation $\Lambda$ with ${\cal C}(\Lambda,\mathbb{P})>0$ can be exactly implemented with a finite amount of coherence cost, as the lower bound diverges as $\epsilon$ approaches 0. Therefore, the problem of whether cost-diverging Gibbs-preserving Operations exists reduces to whether there exists a pairwise reversible Gibbs-preserving Operation $\Lambda$ and a reversible pair $\mathbb{P}$ such that ${\cal C}(\Lambda,\mathbb{P})>0$ at all. The following result not only shows the existence of such operations but provides a continuous family of those.
@@
We prove this in Sec. III of the Supplemental Material, which also provides an explicit construction of the corresponding pairwise reversible Gibbs-preserving Operation. Here is an illustrative example encompassed in Theorem 2. Let $S$ and $S^{\prime}$ be qubit systems with $H_{S}=|{1}\rangle\!\langle{1}|$ and $H_{S^{\prime}}=0$. Since $\tau_{S,0}=1/(1+e^{-\beta})$, $\tau_{S,1}=e^{-\beta}/(1+e^{-\beta})$, and $\tau_{S^{\prime},0}=\tau_{S^{\prime},1}=1/2$, these systems satisfy (9) for arbitrary finite temperature. The corresponding pairwise reversible Gibbs-preserving Operation $\Lambda:S\to S^{\prime}$ is
@@
It is insightful to see the structural difference between the Gibbs-preserving Operations in Eqs. (2) and (10). The one in (2) can create coherence from an incoherent state input state $|{1}\rangle\!\langle{1}|$. On the other hand, the channel in (10) cannot create coherence at all—in fact, output states are always incoherent for any input states. Instead, it can perform a measurement in a coherent basis. As we show in Sec. IV of the Supplemental Material, the coherent cost for the channel in (2) is upper bounded by ${\cal F}(\eta)+{\cal F}(\sigma)$, which corresponds to the sum of coherence that can be created by the channel. This shows a drastic asymmetry between creation and detection of coherence when it comes to its realization.
@@
Theorems 1 and 2 provide an overview of the classification of Gibbs-preserving Operations (Fig. 1). We remark that not all pairwise reversible Gibbs-preserving Operations come with a diverging coherence cost (e.g., identity channel)—Theorem 1 ensures the infinite cost only when there is a reversible pair $\mathbb{P}$ satisfying ${\cal C}(\Lambda,\mathbb{P})>0$. On the other hand, our results do not rule out the possibility that all cost-diverging Gibbs-preserving Operations are pairwise reversible. We also show that among the set of Gibbs-preserving Operations, cost-diverging ones are atypical. Nevertheless, we stress that many well-structured Gibbs-preserving Operations of interest can still come with diverging or extremely high coherence cost—see Sec. V of the Supplemental Material for details.
@@
We prove this in Sec. VI of the Supplemental Material, which also provides an explicit expression of the pair $(\rho,\sigma)$ of states. In the proof, we also show that the approximate realization of the cost-diverging state transformation requires the cost inversely proportional to the error. Theorem 3 shows that when the available coherence resource is restricted to be finite, the range of transformations achievable by Gibbs-preserving Operations is strictly narrowed. We stress that the condition (9) is extremely mild—for virtually all scenarios Theorem 3 puts nontrivial restrictions on the physically available state transformations by Gibbs-preserving Operations.
@@
Proof can be found in Sec. VII of the Supplemental Material. This particularly ensures that the quantity ${\cal C}(\Lambda,\mathbb{P})$ defined in (6) serves as a key quantity that characterizes the coherence cost for a certain class of Gibbs-preserving Operations, which in turn provides an operational interpretation to this quantity.
Theorem 4 ensures the existence of a Gibbs-preserving Operation that almost achieves the lower bound. However, it does not tell much about the general upper bound that could be applied to an arbitrary Gibbs-preserving Operation. In the following, we show that, by giving up obtaining the form that almost matches the lower bound, we can obtain the general sufficient coherence cost that can be universally applied to all Gibbs-preserving Operations. In fact, we find that the applicability of our bound is much beyond Gibbs-preserving Operations—it gives a sufficient coherence cost for an arbitrary quantum channel. (Proof in Sec. VII of the Supplemental Material.)
@@
for some environments $E$ and $E^{\prime}$ such that $S\otimes E=S^{\prime}\otimes E^{\prime}$, some unitary $V$ on $S\otimes E$, and some pure incoherent state (i.e. an energy eigenstate) $|\eta\rangle$ on $E$. Then,
@@
Let us now remark a unique characteristic of our results in relation to the previous result for trivial Hamiltonian. When input and output states are only equipped with trivial Hamiltonian, every state becomes an incoherent state, i.e., invariant under Hamiltonian evolution, and thus coherence loses its status as a precious resource. Therefore, the meaningful question in such a setting is to ask the work cost (the minimum number of work bit required) for implementing unital channels using Noisy Operations Horodecki et al. (2003), which respectively corresponds to the Gibbs-preserving and Thermal Operations for trivial Hamiltonian. Ref. Faist et al. (2015b) showed that the minimum work cost is upper bounded by a quantity scaling as $\sim\log(1/\epsilon)$ with implementation error $\epsilon$, which particularly diverges at the limit of exact implementation. Nevertheless, its lower bound has still not been established, and therefore it is still unclear if this diverging cost is a fundamental phenomenon or merely an artifact of their specific construction, which is based on a decoupling technique Rio et al. (2011). On the other hand, our Theorem 1 provides a lower bound that scale with $1/\epsilon$, which is complemented by upper bounds in Theorems 4 and 5 with the same scaling. To the best of our knowledge, our results are the first ones that establish the inherently diverging thermodynamic cost for implementing Gibbs-preserving Operations.
@@
Acknowledgements.
We thank Kaito Watanabe and Bartosz Regula for fruitful discussions. H.T. was supported by JSPS Grants-in-Aid for Scientific Research No. JP19K14610, No. JP22H05250, and No. JP25K00924, and MEXT KAKENHI Grant-in-Aid for Transformative Research Areas B “Quantum Energy Innovation” Grant Numbers 24H00830 and 24H00831, JST PRESTO No. JPMJPR2014, JST MOONSHOT No. JPMJMS2061, and JST FOREST No. JPMJFR2365. R.T. acknowledges the support of JSPS KAKENHI Grant Number JP23K19028, JP24K16975, JP25K00924, JST, CREST Grant Number JPMJCR23I3, Japan, and MEXT KAKENHI Grant-in-Aid for Transformative Research Areas A “Extreme Universe” Grant Number JP24H00943.
Supplemental Material
I Background and setting
I.1 Quantum resource theories
@@
$\Lambda(\sigma)\in\mathbb{F},\enskip\forall\Lambda\in\mathbb{O},\enskip{\color[rgb]{0,0,0}\forall\sigma\in\mathbb{F},}$ (S.14)
for a set $\mathbb{F}$ of free states and a set $\mathbb{O}$ of free operations. Employing the flexibility in choosing different sets of free operations is usually effective when $\mathbb{O}$ comes with a complicated structure and is difficult to analyze. For instance, in the case of entanglement, it is notoriously hard to study the full potential of LOCC, and therefore several classes of other operations, which include LOCC as their subset, were investigated. One such set is separability-preserving operations, which is the maximal set that satisfies (S.14). This admits a great simplification of the analysis and results in significant insights into entanglement transformation Brandão and Plenio (2010); Brandao and Datta (2011); Regula et al. (2019).
@@
I.2 Thermal and Gibbs-preserving Operations
@@
$\Lambda(\rho)=\mathrm{Tr}_{E^{\prime}}\left(U\rho\otimes\tau_{S}U^{\dagger}\right),\quad[U,H_{\rm tot}]=0$ (S.16)
where $H_{\rm tot}=H_{S}\otimes\mathds{1}_{E}+\mathds{1}_{S}\otimes H_{E}=H_{S^{\prime}}\otimes\mathds{1}_{E^{\prime}}+\mathds{1}_{S^{\prime}}\otimes H_{E^{\prime}}$ is the total Hamiltonian. In this manuscript, we call a closure of the set of Thermal Operations simply Thermal Operations, to which all of our results equally apply. On the other hand, one can also consider a broader class based on the axiomatic formulation, which is only restricted by the minimum requirement that they should map Gibbs states to Gibbs states. This is known as Gibbs-preserving Operations—a channel $\Lambda:S\to S^{\prime}$ is called a Gibbs-preserving operation if
$\Lambda(\tau_{S})=\tau_{S^{\prime}}.$ (S.17)
@@
I.3 Quantum coherence in energy eigenbasis
@@
$\Pi_{E}=\sum_{n:E_{n}=E}|{n}\rangle\!\langle{n}|.$ (S.18)
Then, a state $\rho$ has energetic coherence if $\sum_{E}\Pi_{E}\rho\Pi_{E}\neq\rho$. Equivalently, a state $\rho$ has nonzero coherence if there exists $t\in\mathbb{R}$ such that $e^{-Ht}\rho e^{iHt}\neq\rho$.
The letter expression particularly allows us to formalize energetic coherence in relation to a group action represented by a unitary representation of ${\rm U}(1)$ (or $\mathbb{R}$ if the Hamiltonian contains relatively irrational eigenvalues). Namely, incoherent states—states that do not have coherence—are equivalent to the states invariant under a unitary representation $\{e^{-iHt}\}_{t}$. This observation provides a way of quantifying the amount of energetic coherence employing the resource theory of asymmetry Gour and Spekkens (2008); Marvian Mashhad (2012); Zhang et al. (2017); Takagi (2019); Kudo and Tajima (2023); Yamaguchi and Tajima (2023a, b); Shitara and Tajima (2023), which considers states invariant under action of a unitary representation of a group $G$ as free states, i.e., $\mathbb{F}=\{\sigma\,|\,U_{g}\sigma U_{g}^{\dagger}=\sigma,\ \forall g\in G\}$, and the operations covariant with such group actions as free operations, i.e., $\mathbb{O}=\{\Lambda:S\to S^{\prime}\,|\,U_{S^{\prime},g}\Lambda(\rho)U_{S^{\prime},g}^{\dagger}=\Lambda(U_{S,g}\rho U_{S,g}^{\dagger}),\ \forall g\in G\}$ where $U_{X,g}$ is a unitary representation acting on a system $X$. One can then realize that incoherent states coincide with the set of free states in the framework of resource theory of asymmetry with U(1) group with the representation $U_{t}=e^{-iHt}$, equipped with
$\mathbb{F}_{\rm inc}=\{\sigma\,|\,e^{-iHt}\sigma e^{iHt}=\sigma,\ \forall t\}$ (S.19)
@@
$\mathbb{O}_{\rm cov}=\{\Lambda:S\to S^{\prime}\,|\,e^{-iH_{S^{\prime}}t}\Lambda(\rho)e^{iH_{S^{\prime}}t}=\Lambda(e^{-iH_{S}t}\rho e^{iH_{S}t})\ \forall t\}.$ (S.20)
@@
$\Lambda(\rho)=\mathrm{Tr}_{E^{\prime}}\left(U\rho\otimes\sigma U^{\dagger}\right),\quad[U,H_{S}\otimes\mathds{1}_{E}+\mathds{1}_{S}\otimes H_{E}]=0,\ \sigma\in\mathbb{F}_{\rm inc}$ (S.21)
for some systems $E$ and $E^{\prime}$. The similarity between (S.16) and (S.21) represents an intriguing interplay between resource theories for quantum thermodynamics and energetic coherence. Indeed, since Gibbs states do not have energetic coherence by definition, we immediately notice that
$\mathbb{O}_{\rm TO}\subseteq\mathbb{O}_{\rm cov}.$ (S.22)
@@
I.4 Coherence cost for Gibbs-preserving Operations